OpenStudy (anonymous):

Find the future value of an annuity if you invest $850 quarterly for 5 years at 1.5% compounded quarterly.

OpenStudy (tkhunny):

Basic Principles. Define the building blocks and write ti down. i = 0.015 -- The annual percentage rate. It is given in the problem statement. j = i/4 = 0.00375 -- The quarterly percentage rate r = 1+j = 1.00375 -- The Quarterly Accumulation Factor. $1 invested now will be $1 * r = $1.015 in 3 months. Now build. \( 850⋅(r^{20}+r^{19}+r^{18}+...+r)=850⋅\dfrac{r^{21}−r}{j}\) No need to wonder which formula to use. Define what you need and BUILD IT!

OpenStudy (wolf1728):

Interest = (1 + .015/4)^4 Interest = 1.00375^4 Interest = 1.0150845861 Interest = 1.50845861% Annuity Value Here's the Annuity Formula

OpenStudy (wolf1728):

We know that $3,400 is being invested per year so 3,400 * [((1.0150845861)^6 -1) / .0150845861] - 3,400 3,400 * (.0939901174 /.0150845861) -3,400 3,400 * (6.2308714854) -3,400 21,184.96 -3,400 = 17,784.96 See? It's just that simple!

OpenStudy (wolf1728):

After all that, it is still not an exact answer. The formula I posted (and used) is for an annuity which has amounts added annually. I should have used a quarterly annuity formula - but I couldn't locate one. Anyway, that answer should be close enough.

OpenStudy (tkhunny):

$17,685.54 That is as exact as possibly to the penny and is based on the exact quarterly formula given above.